When 1.1.2g of iron reacts with 32g of bromine , iron (III)bromide is produced. Find the mass of iron (III) bromide.

1
Expert's answer
2022-05-17T06:28:03-0400

2Fe + 3Br2 = 2FeBr3


n(Fe) = 11.2 g / 55.8 g/mol / 2 = 0.1 moles

n(Br2) = 32 g / 159.8 g/mol / 3 = 0.07 moles


Fe - excess reagent

Br2 - limiting reagent


m(FeBr3) = (m(Br2) * 2Mr(FeBr3)) / 3Mr(Br2) = (32 g * 2*295,56 g/mol) / 3*159.8 g/mol = 39.46 g